OpenStudy (anonymous):

What is the solution to the equation 1/3(9x - 6) = 2x + 3? x =

OpenStudy (hockeychick23):

distribute on the left side first, do you know how to do that?

OpenStudy (anonymous):

not really

OpenStudy (igreen):

First distribute 1/3 into the parenthesis by multiplying 1/3 to each term inside of it. \(\dfrac{1}{3} \times 9x =~?\) \(\dfrac{1}{3} \times -6 =~?\) Can you multiply those?

OpenStudy (hockeychick23):

ok so you need to multiply 9x and -6 by 1/3, so first what is 1/3(9x) ?

OpenStudy (anonymous):

No it's 9x minus 6...sorry

OpenStudy (hockeychick23):

yep, but based on the distributive property, you need to multiply both 9x and -6 by 1/3

OpenStudy (igreen):

^^

OpenStudy (anonymous):

so we turn it into a negative?

OpenStudy (igreen):

So can you multiply those two things I just gave you?

OpenStudy (hockeychick23):

yes, adding a negative and subtracting are the same thing

OpenStudy (anonymous):

Okay I don't know this stuff sorry, but yeah I can multiply them

OpenStudy (igreen):

Okay, tell me what you get.

OpenStudy (anonymous):

i got 3 for the first one and -2 for the second one

OpenStudy (hockeychick23):

yep thats correct, so your new equation is 3x-2=2x+3, do you know what to do next?

OpenStudy (anonymous):

Not really

OpenStudy (igreen):

Yep, we have: \(3x - 2 = 2x + 3\) Now add 2 to both sides.

OpenStudy (hockeychick23):

ok so you want to get all the x's on one side and all the numbers on the other, so subtract 2x from both sides and tell me what you get

OpenStudy (igreen):

Tell me what you come up with

OpenStudy (anonymous):

Is this from a test ? @TFIOS_GUS623

OpenStudy (anonymous):

It's from an assignment @101

OpenStudy (anonymous):

Is the answer -1 and 3 or is it 3 and 7? I'm confused...

OpenStudy (hockeychick23):

ok so your equation is : 3x-2=2x+3, subtract 2x from both sides

OpenStudy (anonymous):

okay so -1 for the first and 3 for the second?

OpenStudy (hockeychick23):

no, you have to use the equation i gave you, 3x-2-2x=2x-2x+3 this= x-2=3, now add 2 to both sides

OpenStudy (anonymous):

I'm just....very confused right now

OpenStudy (hockeychick23):

ok what minus 2 gives you 3?

OpenStudy (anonymous):

@TFIOS_GUS623 Questoins from tests are not allowed on here.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

*Questions

OpenStudy (hockeychick23):

correct, 5 is your answer

OpenStudy (anonymous):

Like I said, it's NOT from a TEST. @101

OpenStudy (anonymous):

So, 5 is my overall answer?

OpenStudy (anonymous):

Then why are you posting these questions one after another when someone helps you ?

OpenStudy (anonymous):

It seems you are copy and pasting them, how else can you post them so quickly.

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

Like I said before...it's an ASSIGNMENT. Just homework. It's not a test.

OpenStudy (hockeychick23):

yes 5 is your overall answer, you can plug it in and check: 3(5)-2=2(5)+3 --> 13=13

OpenStudy (anonymous):

Thank you @hockeychick23

OpenStudy (uri):

@101 He's trying to learn stop annoying him.

OpenStudy (hockeychick23):

no problem (:

OpenStudy (igreen):

@uri Ikr?

OpenStudy (anonymous):

That is quite hard to believe. @TFIOS_GUS623 I myself come from an online school and I know what assignments are. They are nothing like these questions.

OpenStudy (anonymous):

I would like to say something. We go to different online schools. This is from a five question assignment. It is NOT a test, I don't have a scheduled test until after my Christmas break. I would like to say I do not appreciate you reprimanding me for needing help on a website you go to for HELP. My business is my business and frankly you interrogating me is uncalled for.

OpenStudy (anonymous):

@101 So thank you and have a nice day.